# Direct observation of the crystal electric-field splitting under magnetic field and uncovering field-induced magnetic phase transition in triangular rare-earth magnet CsErSe2

### v1

@@ V Summary
Neutron Refinments: We refined the magnetic structure of CsErSe2 by performing irreducible representational analysis for the $P6_{3}/mmc$ space group and the $(1/2,0,1)$ propagation vector. The four irreducible representations and further information is available in Sup . There are many possible combinations of irreducible representations that will account for this magnetic structure, but we here show the simplest option that preserves uniform moment size: alternating canting along the $c$-axis, shown in Fig. 8.
@@
Figure 9: (a) Convergence of phonons with respect to cell size, using a $4\times 4\times 1$ (128 atom) supercell of the primitive cell.
In Fig. 9 we show phonon dispersions for $4\times 4\times 1$ (128 atom) supercells. The close agreement shows that either supercell size could be used, though we use the $4\times 4\times 1$ cell for all final results. Note that longitudinal orbital – transverse orbital (LO-TO) splitting is included in these calculations through the non-analytic term correction (NAC) as implemented in Phonopy. Table 3 listed the $\Gamma$ point energy from Fig. 9 that were used to compare the field-independent Raman shift in Fig. 6. Further information on the calculations and the convergences are available in Sup .
mode # modes Energy (cm-1)
@@
Table 3: $\Gamma$ point mode energy in cm-1 as shown in Fig. 9. Extended version of this table is available in Sup .

### v2

@@ Acknowledgements.
Neutron Refinements: We refined the magnetic structure of CsErSe2 by performing irreducible representational analysis for the $P6_{3}/mmc$ space group and the $(1/2,0,1)$ propagation vector. The four irreducible representations and further information is available in Sup . There are many possible combinations of irreducible representations that will account for this magnetic structure, but we here show the simplest option that preserves uniform moment size: alternating canting along the $c$-axis, shown in Fig. 8.
@@
Figure 9: Phonon dispersions calculated using density functional theory with a $4\times 4\times 1$ supercell.
In Fig. 9 we show phonon dispersions for $4\times 4\times 1$ (128 atom) supercells. The close agreement shows that either supercell size could be used, though we use the $4\times 4\times 1$ cell for all final results. Note that longitudinal orbital – transverse orbital (LO-TO) splitting is included in these calculations through the non-analytic term correction Pick et al. (1970) as implemented in Phonopy. Table 3 listed the $\Gamma$ point energy from Fig. 9 that were used to compare the field-independent Raman shift in Fig. 6. Further information on the calculations and the convergences are available in Sup .
mode # Mulliken symbol Energy (cm-1)
@@
Table 3: Calculated $\Gamma$ point energy of phonon modes in Mulliken symbol as shown in Fig. 9. Extended version of this table is available in Sup .